DX 4D flattening of TS3 - 20180208

20:00 Tertiary in postion. Ready to search

20:42 Turn off LUCI2 cold head #1 and treehouse ventilation fans (LUCI1 off telescope); LUCI2 cold head #2 still on

20:45 Put act 376 and 555 into no-pos list.
  • forces now -0.07 to 0.004.N
  • before -0.31 to 0.34 N
  • hex 0.583, -0.864, 3.055, 215.65, 165.84
  • images from last night at ~54 elevation
  • Actuators with high force:
    high_neg_force_376_216_104.jpg
  • Actuators with high force:
    high_pos_force_555.jpg

21:00 Switch over from summer chiller to winter chiller testing

21:05 Flatten with 20180205_225511
  • flatten with 50
    • flat

21:30 align pupil by moving Tertiary
  • move coma free pointing along with tertiary tip/tilt
    • Coma Free Tip +1" == -9" tip, -9" tilt
    • Coma Free Tilt +5" == -33" tip, 55" tilt

21:36 Winter chiller back on after being off for awhile.

21:54 Turn off LUCI2 cold head #2 (to get rid of 18 Hz peaks)

22:25
  • Hex 2.111, -0.2612, 3.058, 72.159, 449.35
  • Tertiary -479.9, -208, 0, 56.41
  • flatten 30 modes = 820 nm rms
  • flatten 200 modes = 102 nm rms flat 223318
  • push with gaussian at 470, 471, 478 91 nm rms

22:34 Noise data
  • 223400_noise
  • 223500_noise

22:54 Push 91 nm
  • 20180208_225400_300mode_with_push.sav

  • print, i4d_opt_intmat(0, 200)
    • 225814 (3342, 3160 nm)
  • print, i4d_opt_intmat(201, 400)
    • 225927 (3530, 3333 nm)
  • print, i4d_opt_intmat(401, 600)
    • 230050 (3760, 3414 nm)

23:01 Noise
  • 20180208_230100_noise (47 us)
  • 21080208_230200_noise (31 us)
  • 21080208_230300_noise (31 us)
  • 21080208_230400_noise (31 us)

23:05 Hand over to mountain
  • hex 2.111, -0.2612, 3.058, 72.15, 449.37
  • tertiary -479, -207 0 56.4

tn1=['20180208_223400_noise','20180208_223500_noise','20180208_230100_noise']
tn2=['20180208_230200_noise','20180208_230300_noise','20180208_230400_noise']

-- %USERSIG{DougMiller - 2018-02-08}%

Summary

I think we've made some progress on the vibration front. On Monday, we demonstrated that the LUCI2 cold head #1 was making all of the harmonic noise seen in the OVMS TT PSD in the region of 60 Hz (all are harmonics of the 1.2 Hz cold head stroke rate). All the other three cold heads contribute close to nothing at 60 Hz.

Today, we turned off both LUCI2 coldhead #1 (LUCI1 off telescope) and the 4 treehouse ventilation fans (20:42 UT), but we still had a large vibration peak near 18 Hz. When we turned off LUCI2 coldhead #2 (21:54 UT), the peak at 18 Hz went away. So it seems that LUCI2 coldhead #2 has significant harmonics at 18 Hz, but not at 60 Hz, and LUCI2 cold head #1 is vice versa. When we turn off treehouse fans and LUCI coldheads, the vibration power in the 8-20 Hz window goes down by a factor of roughly 2 as measured by OVMS. Doug is now trying to get 4D data that demonstrates this improvement.

Now that we've removed the strong peaks, the vibration power is dominated by the broad eigenmodes of the swing arms. I've not yet understood what is pumping those eigenmodes.

THERE IS NO 4D DATA IN THESE PLOTS. THEY ARE ONLY FROM OVMSplus.

The upper panel is the X=Tilt time series of focal plane image motion predicted by OVMSplus based on the OVMS accelerometers. Black is SX, Blue is DX.

The second panel is the PSD of the above time series from 0.1 to 84 Hz.

The large-ish black numbers are the overall rms for SX and DX respectively. The large-ish red numbers are corrected to only those frequencies that 4D is most sensitive to (mostly removing stuff below 7 Hz).

The third and fourth panels are the same for the Y=Tip direction.

One plot has LUCI coldheads and treehouse fans on, and the other has them off (as discussed above). Note in the "Off" plot, that the Red DX X rms has decreased from 11 mas to 2.4 mas rms.

  • PSD from OVMSplus with Treehouse Fans and LUCI cold heads OFF:
    ovms.mode2.20180208LUCIsoffTreehousesoff.jpg

  • PSD from OVMSplus with Treehouse Fans and LUCI cold heads ON:
    ovms.mode2.20180208LUCI2onTreehouseson.jpg
I Attachment Action Size Date Who Comment
high_neg_force_376_216_104.jpgjpg high_neg_force_376_216_104.jpg manage 142 K 09 Feb 2018 - 12:36 DougMiller Actuators with high force
high_pos_force_555.jpgjpg high_pos_force_555.jpg manage 101 K 09 Feb 2018 - 12:36 DougMiller Actuators with high force
ovms.mode2.20180208LUCI2onTreehouseson.jpgjpg ovms.mode2.20180208LUCI2onTreehouseson.jpg manage 1 MB 09 Feb 2018 - 17:11 JohnHill PSD from OVMSplus
ovms.mode2.20180208LUCIsoffTreehousesoff.jpgjpg ovms.mode2.20180208LUCIsoffTreehousesoff.jpg manage 1 MB 09 Feb 2018 - 17:10 JohnHill PSD from OVMSplus
Topic revision: r5 - 09 Feb 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback